how to convert 1.5 m/s2 to 19440km/h^2 step by step..

Grade:12th pass

1 Answers

Khimraj
5 years ago
1m = (1/1000) Km
and 1s = (1/60)m = (1/3600)h
So 1s2 = (1/3600*3600)h2 = (1/12960000)h2
Then 1.5 m/s2 = 1.5*12960000/1000 = 19440km/h2.
Hope it clears. If you like answer then please approve it.